OpenStudy (cruffo):

Find the fixed point(s) of \[f(x) = -2x(x-1).\]

OpenStudy (anonymous):

may be x=0.1

OpenStudy (cruffo):

how did you get that?

OpenStudy (anonymous):

i think u need when f(x)=0

OpenStudy (watchmath):

solve f(x)=x

OpenStudy (cruffo):

The definition of fixed point being used here is a point that does not change on repeated compositions. Or as watchmath said f(x) = x :)

OpenStudy (watchmath):

-2x(x-1)=x x(-2x+2)=x x(-2x+1)=0 x=0 or x=1/2.

OpenStudy (cruffo):

Great!

OpenStudy (anonymous):

watchmath can u tell me why we are doing f(x)=x

OpenStudy (watchmath):

That is the definition of fixed point. c is a fixed point if f(c)=c.

OpenStudy (anonymous):

ok...thanks..